Formula between electric field, mass, charge  please help me out l am very nervous 

Grade:12

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

F = k q1*q2/ r^2

E = F/q1

Now F= m*a or m*v/t

E = mv/tq

Regards